A company currently pays a dividend of $4 per share (D0= $4). It is estimated that the company’s dividend will grow at a rate of 10% per year for the next 2 years, and then at a constant rate of 5% thereafter. The company’s stock has a beta of 1.6, the risk-free rate is 4% and the market risk premium is 2%. What is your estimate of the stock’s current price?

Please show solution in Excel. Thank you!
Answer #1

First we need to calculate the rate of return on equity

Using the CAPM model,

r = Rf + β(ERm - Rf)

where,

r = return on equity
Rf = Risk-free rate = 4%
β = Beta of the stock = 1.6
ERm - Rf = market risk premium = 2%

=> r = 4 + 1.6*2 = 7.2%

Calculate in Excel as below -

B7 2 CAPM model 4 Risk Free Rate 5 Beta 6 Market Risk Premium 7 Return on Equity 4% 1.6 2% 7 20%

Let current price of stock be P0

Let current period be Period 0, and Dividend in Period 0 = D0

Given D0 = 4

D1 = 4*1.10 = 4.40

D2 = 4.40*1.10 = 4.84

Using Gordon Growth model,

Hence, P1 = D2/(r-g) = 4.84/(0.072-0.05) = 220

P0 = P1/(1+r) = 220/1.072 = $205.22

Calculate in excel -

G4 fz -F5/(B7-596) 2 CAPM model Dividend Price of Stock Formula Year 0 Year 1 Year 2 205.22-G4/(1+87) 2201-F5/B7-5%) 4 Risk Free Rate 5 Beta 6 Market Risk Premium 7 Return on Equity 4% 1.6 2% 7.20% 4.4 4.84